OpenStudy (anonymous):

A firecracker shoots up from a hill 140 feet high with an initial speed of 100 feet per second. Using the formula H(t) = −16t2 + vt + s, approximately how long will it take the firecracker to hit the ground?

OpenStudy (zzr0ck3r):

-16t^2+100t+140=0 make sense?

OpenStudy (anonymous):

I understand that part it's after that that confuses me

OpenStudy (zzr0ck3r):

do you know the quadratic formula?

OpenStudy (zzr0ck3r):

-16t^2 + 100t+140 = 0 -8t^2+50t+70=0 -4t^2+25t+35=0 t^2-25/4t-35/4=0 t^2-25/4t = +35/4 (t-25/8)^2 = 35/4+(25/8)^2 t-25/8 = +-sqrt(35/4+(25/8)^2) t = +-sqrt(35/4+(25/8)^2)+25/8

OpenStudy (anonymous):

Yes, but everytime I solve one of these equations that says " aprroximately how long does this object take to hit the ground" I always mess up. Idk why

OpenStudy (zzr0ck3r):

this is the method of completing the square. Do you know it?

OpenStudy (zzr0ck3r):

or solve this -100+-sqrt(100^2-4(-16)(140))/(2(-16))

OpenStudy (anonymous):

No. But when youtry to show me is there any way you could draw it out because you typing it just confuses me even more.

OpenStudy (zzr0ck3r):

drawing on here is very hard, I cant really draw out equations.

OpenStudy (anonymous):

oh sorry I've never tried it

OpenStudy (anonymous):

And I did the last equation you put up and got an answer in the hundreds?